OpenStudy (anonymous):

I AWARD MEDALS!!! Can someone please show me the steps to solving this?? I have no clue Each level of a smartphone app adds more gems for you to match. On level one, there were 13 gems. On level twelve, there were 123 gems. When you complete all twelve levels, how many total gems will you have matched?

OpenStudy (gorv):

level 1= 13 level 12= 123 total gems added=123-13=110 in 11 levels(2nd 3rd 4th 5th 6th 7th 8th 9th 10th 11th 12th level ...total 11 level) so gems added on each level= 110/11=10 gems on each level

OpenStudy (gorv):

so add gems of each level ist=13 2nd=13+10 3rd=13+10 +10 ....... . . . . . . . 11th= gems on 10th level +10 12th=123 add all of them=matchedgems

OpenStudy (anonymous):

I got 803 but that isn't an option

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

13+23+33+43+53+63+73+83+93+103+113+123=816

OpenStudy (gorv):

i think u forgot to add ist level=13

OpenStudy (anonymous):

Oh! Yeah. sorry. Thank you!